# Extremal metrics involving scalar curvature

### v2

@@ Abstract
We investigate extremal metrics at which various types of rigidity theorems involving scalar curvatures hold. The rigidity we discuss here is related to the rigidity theorems presented by Mario Listing in his previous preprint. More specifically, we give some sufficient conditions for metrics not to be rigid in this sense. We also give several examples of Riemannian manifolds that satisfy such sufficient conditions.
@@ 1 Introduction
Llarull [21] showed some rigidity results for the standard sphere. And, Goette and Semmelmann [10] generalized it to locally symmetric spaces of compact type and nontrivial Euler characteristic. Later Listing [19] generalized their results in the following form.
Theorem 1.1 ([19, Theorem 1]).
Let $(M_{0}^{n},g_{0})~{}(n\geq 3)$ be an oriented closed Riemannian manifold with nonnegative curvature operator, positive Ricci curvature and non-vanishing Euler characteristic. Suppose that $(M^{n},g)$ is an oriented closed Riemannian manifold and $f:M\rightarrow M_{0}$ is a spin map of non-zero degree. If the scalar curvature satisfies
@@
For the case that $M=M_{0}$ and $f=\mathrm{id}_{M}$, he also gave the following type of rigidity theorem.
Theorem 1.2 ([19, Theorem 2]).
Suppose $(M^{n},g_{0})$ is an oriented Riemannian manifold of dimension $n=4k+1~{}(k\in\mathbb{N})$ with nonnegative curvature operator, positive Ricci curvature and non-vanishing Kervaire semicharacteristic $\sigma(M)\neq 0$. If $g$ is a Riemannian metric on $M$ satisfying
@@
and a functional $\mu_{g_{0}}$ on $\mathcal{M}$, which is determined by $g_{0}$ and scaling invariant of weight $-1$, i.e., $\mu(c\cdot g)=c^{-1}\mu(g)$ for all $c>0$ and $g\in\mathcal{M}$. If the metric $g_{0}$ is rigid with respect to the functional $\mu_{g_{0}}$ in a certain sense, then $\mu_{g_{0}}$ is an upper bound of $R_{min}$ as a functional on $\mathcal{M}$ and these values coincide at $g_{0}$. On the other hand, when $M$ is closed $n$-manifold with non-positive Yamabe invariant $Y(M)$, the following functional is an upper bound of $R_{min}$ on $\mathcal{M}$:
@@
Moreover, if a smooth metric $g_{0}$ attains the equality, then it is a Yamabe metric (i.e., $Y(M,[g_{0}])=Y(M)$) and an Einstein metric. However, when the Yamabe invariant is positive, this is not the case in general (see Remark 2.1 below or [20, Chapter 3]). A Yamabe metric is expected to be standard in some sense (cf. [28, Section 1]), but it remains a difficult problem to know how to actually obtain it as a limit of the sequence of solutions to the Yamabe problem, and in what sense it is standard (see also Subsection 6.2 below).
@@
    $\begin{split}&~{}~{}~{}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n-1}>0\\ &\left(\mathrm{resp.}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,2}>0\right)\end{split}$
@@
    $\begin{split}&~{}~{}~{}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n}>0\\ &\left(\mathrm{resp.}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,1}>0\right)\end{split}$
@@
  • All the assumptions of the above Theorems 1.3, 1.4, 1.5 and 1.6 especially imply that
@@
Interestingly, Dahl–Kröncke [7, 16] recently also discovered a relation between stability of Einstein metrics and certain type of scalar curvature rigidity. Since every Einstein metric does not satisfy any of the above assumptions (see Remark 1.2 above), our theorems above cannot be applied to Einstein metrics. However, we can also prove the following.
Theorem 1.7.
Let $(M^{n},g)$ be a closed non-positive Einstein manifold of dimension $n\geq 2$ with unit volume. Assume either that
  • $(M,g)$ is Ricci flat, unstable ([16]), i.e., there is a tt-tensor $h$ (see 4 for the definition) such that
[…]
    where $\Delta_{E}:=\nabla_{g}^{*}\nabla_{g}-2\overset{\circ}{R}$ is the Einstein tensor and
    $(\overset{\circ}{R}h)(X,Y):=\sum_{i}h(R(X,e_{i})Y,e_{i}).$
    Moreover, there is a corresponding curve of metrics $g_{s}~{}(|s|<\varepsilon)$ lying in the space of all Riemannian metrics of unit volume with constant scalar curvature such that $\left.\frac{d}{ds}g_{s}\right|_{s=0}=h$ and $\left.\frac{d^{2}}{d^{2}s}g_{s}\right|_{s=0}$ is also a tt-tensor,
  • $R_{g}<0$ and there is a non-vanishing tt-tensor $h$, i.e.,
    $\max_{M}\sup\{C\in\mathbb{R}\mid~{}h_{x}(v,v)\geq C\cdot g_{x}(v,v)~{}\mathrm{for~{}all}~{}v\in T_{x}M\}<0.$
Then, there is a curve of metrics $g_{s}$ defined for sufficiently small $s>0$ with $\frac{d}{ds}g_{s}|_{s=0}=h$ such that $g_{s}$ satisfies
$R_{g_{s}}>R_{g}\cdot\|g\|_{1,g_{s}}^{2}$ (10)
on $M$. Here, $\|g\|_{1,g_{s}}$ is the function defined in (3).
Remark 1.3.
All known examples of unstable Einstein manifolds are of positive scalar curvature (see [15]). Therefore, it is not known whether there is a metric that satisfies the first condition.
Question 1.1 ([8, p. 65]).
Are all compact Einstein manifolds with non-positive scalar curvature stable?
Theorem 1.8.
Let $(M^{n},g)$ be a closed Einstein manifold of dimension $n\geq 2$ with unit volume. Assume that $(M,g)$ is not isometric to the standard round sphere. For a tt-tensor $h$ (see 4 for the definition), we denote the second eigenvalue of $h$ with respect to $g$ as $\lambda^{h}_{2}$. Assume either that
  • $R_{g}\neq 0$ and there is a tt-tensor $h$ such that $\max_{M}\lambda^{h}_{2}<0$, or
  • $R_{g}=0$, $g$ is unstable and a tt-tensor $h$ satisfying (9). Moreover, there is a corresponding curve of metrics $g_{s}~{}(|s|<\varepsilon)$ lying in the space of all Riemannian metrics of unit volume with constant scalar curvature such that $\left.\frac{d}{ds}g_{s}\right|_{s=0}=h$ and $\left.\frac{d^{2}}{d^{2}s}g_{s}\right|_{s=0}$ is also a tt-tensor.
Then, there is a curve of metrics $g_{s}~{}(|s|<\varepsilon)$ such that for all $s\in(0,\varepsilon)$, $g_{s}$ satisfies
$R_{g_{s}}>R_{g}\cdot\|g\|_{2,g_{s}}$
on $M$. Here, $\|g\|_{2,g_{s}}$ is the function defined in (6).
Remark 1.4.
In the case of $R_{g}\neq 0$, the assumption “$\max_{M}\lambda^{h}_{2}<0$” especially implies that $h$ is non-vanishing on $M$.
According to Listing’s work [19], we define three types of rigidities of metrics involving scalar curvature.
@@
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type I scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type II scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type III scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type IV scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
    $\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}-R_{g}(x)\cdot\lambda_{Ric,n-1}(x)\leq 0$
@@
Corollary 1.5.
Let $(M^{n},g)$ be a closed negative Einstein manifold of dimension $n\geq 2$ with unit volume. If it is type I scalar curvature rigid in the sense of Listing, then $h_{x}=0$ at some point $x\in M$ for any tt-tensor $h$.
Corollary 1.6.
Let $(M^{n},g)$ be a closed non-Ricci-flat Einstein manifold of dimension $n\geq 2$ with unit volume that is not isometric to the standard sphere. If it is type II scalar curvature rigid in the sense of Listing, then $\min_{M}\lambda^{h}_{2}\geq 0$ for any tt-tensor $h$. Here, $\lambda^{h}_{2}$ denote the second eigenvalue of $h$ with respect to $g$.
@@ 2 Previous rigidity results for metrics with positive scalar curvature
Let $(M_{0}^{n},g_{0})~{}(n\geq 3)$ be an oriented closed Riemannian manifold with nonnegative curvature operator, positive Ricci curvature and non-vanishing Euler characteristic. Suppose that $(M^{n},g)$ is an oriented closed Riemannian manifold and $f:M\rightarrow M_{0}$ is a spin map of non-vanishing $\hat{A}$-degree $\mathrm{deg}_{\hat{A}}(f)\neq 0$ and $\mathrm{area}(f)\leq 1$. Then $R_{g}\geq R_{g_{0}}\circ f$ implies that $R_{g}=R_{g_{0}}\circ f$. If moreover, $\mathrm{Ric}_{g}>0$ and $02\,\mathrm{Ric}_{g}-R_{g}\cdot g<$ 0 on $M$, then $f:M\rightarrow M_{0}$ is a Riemannian submersion.
@@
Let $(M,g)$ be a compact connected oriented Riemanniam manifold with nonnegatuve curvature operator on $\Lambda^{2}TM$, such that the universal covering of $M$ is homeomorphic to a symmetric space $G/K$ of compact type with $\mathrm{rk}G\leq\mathrm{rk}K+1$. Then $g$ is (globally) area-extremal. If moreover, $\mathrm{Ric}_{g}>0$ and $02\,\mathrm{Ric}_{g}-R_{g}\cdot g<$ 0 on $M$, then $R_{h}\geq R_{g}$ and $h\geq g$ on $\Lambda^{2}TM$ implies $h=g$.
@@
For a fixed Riemannian metric $g_{0}\in\mathcal{M}(M)$, we define the following two functionals $F_{1}(g_{0})$ and $F_{2}(g_{0})$.
$\begin{split}F_{1,g_{0}}&:\mathcal{M}(M)\rightarrow\mathbb{R}~{};~{}g\mapsto R_{g_{0}}\cdot\|g_{0}\|^{2}_{1,g},\\ F_{2,g_{0}}&:\mathcal{M}(M)\rightarrow\mathbb{R}~{};~{}g\mapsto R_{g_{0}}\cdot\|g_{0}\|_{2,g}.\end{split}$
@@
Suppose there is a metric $g\in\mathcal{M}(M)$ such that $R_{min}(g)=\min_{M}R_{g}>F_{1,g_{0}}(g)$. Then, from Definition 1.1, there is a positive constant $c>0$ such that $g=c\cdot g_{0}$. Hence, $R_{g}=c^{-1}R_{g_{0}}=c^{-1}F_{1,g_{0}}(g_{0})=F_{1,g_{0}}(g)$ on $M$. This contradicts our supposition $R_{min}(g)=\min_{M}R_{g}>F_{1,g_{0}}(g)$. The proof for the corresponding statement to $F_{2,g_{0}}$ is similar. ∎
[…]
As pointed out in [20, Section 3], when the Yamabe invariant, alias the sigma constant, $Y(M^{n})$ is non-positive, then
@@ 3 Preliminaries
where the term $g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h$ is a contraction of three copies of $g^{-1}$ (i.e., $g$ with raised indices) and two copies of ¯$\bar{\nabla}h=\bar{\nabla}g$. And, the term $(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}*\mathrm{Ric}_{\bar{g}}$ is the trace of $\mathrm{Ric}_{\bar{g}}$ with respect to $\left((\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}\right)^{-1}$. Note that $\bar{g}+h$ is positive definite if $\|h\|_{\bar{g}}$ is small enough. All the proofs of these formulas are given in Section 7 below.
@@
instead of (17), a more subtle deformation is needed (see [4, Theorem 4]).
@@
$-\Delta_{g}u=\lambda u$
for some positive constant $\lambda>0$. Then, for any $r>0$, there is a small $s>0$ (resp. $s<0$) depending only on $n,g,\lambda,K$ and $r$ such that $g_{s,u}:=(1+su)g$ is a smooth metirc in $B_{r}(p)$ and that
@@
First, we prove the case of $s>0.$ From the formula (16) and Li-Wang’s gradient estimate [18, Theorem 6.1], there is a constant $C>0$ depending on $n$ such that

### v3

@@ Abstract
We investigate extremal metrics at which various types of rigidity theorems involving scalar curvatures hold. The rigidity we discuss here is related to the rigidity theorems presented by Mario Listing in his previous preprint. More specifically, we give some sufficient conditions for the metrics not to be rigid in this sense. We also give several examples of Riemannian manifolds that satisfy such sufficient conditions.
@@ 1 Introduction
Llarull [22] showed some rigidity results for the standard sphere. And Goette and Semmelmann [10] generalized it to locally symmetric spaces of compact type and nontrivial Euler characteristic. Later, Listing [20] generalized their results in the following form.
Theorem 1.1 ([20, Theorem 1]).
Let $(M_{0}^{n},g_{0})~(n\geq 3)$ be an oriented spin closed Riemannian manifold with nonnegative curvature operator, positive Ricci curvature and non-vanishing Euler characteristic. Suppose that $(M^{n},g)$ is an oriented closed Riemannian manifold and $f:M\rightarrow M_{0}$ is a spin map of non-zero degree. If the scalar curvature satisfies
@@
For the case of $M=M_{0}$ and $f=\mathrm{id}_{M}$, he also gave the following type of rigidity theorem.
Theorem 1.2 ([20, Theorem 2]).
Suppose $(M^{n},g_{0})$ is an oriented spin closed Riemannian manifold of dimension $n=4k+1~(k\in\mathbb{N})$ with nonnegative curvature operator, positive Ricci curvature and non-vanishing Kervaire semicharacteristic $\sigma(M)\neq 0$. If $g$ is a Riemannian metric on $M$ satisfying
@@
The condition that a metric has nonnegative curvature operator is preserved under the Ricci flow. Moreover, on a closed manifold, the condition that a metric has positive Ricci curvature is also preserved under the Ricci flow for a sufficiently short time. The Ricci flow solution $g(t)$ is homothetic (i.e. $g(t)=c(t)\phi(t)^{*}g_{0}$ where $c(t)$ is a positive constant and $\phi(t)$ is a diffeomorphism for each $t$) if and only if the initial metric $g_{0}$ is Einstein up to a diffeomorphism. Hence, in dimension three, if $g_{0}$ is non-Einstein metric satisfying the assumption of the above theorem 1.1 or 1.2, then it should be able to obtain a family of metrics (which is a solution of the Ricci flow equation starting at $g_{0}$) that satisfies the assumption of each theorem and is not merely a positive constant multiple of the original metric $g_{0}$. In light of these, we ask the following.
Question 1.1.
Is there any non-Einstein metric $g_{0}$ that satisfies the above Listing’s theorem 1.1 or 1.2?
We will show below that a metric which is not Einstein, and furthermore satisfies a certain assumption, is not a Listing-type extremal metric in a certain sense.
@@
and a functional $\mu_{g_{0}}$ on $\mathcal{M}$, which is determined by $g_{0}$ and the scaling invariant of weight 1-1, i.e., $\mu(c\cdot g)=c^{-1}\mu(g)$ for all $c>0$ and $g\in\mathcal{M}$. If the metric $g_{0}$ is rigid with respect to the functional $\mu_{g_{0}}$ in a certain sense, then $\mu_{g_{0}}$ is an upper bound of $R_{min}$ as a functional on $\mathcal{M}$ and these values coincide at $g_{0}$. On the other hand, when $M$ is closed $n$-manifold with non-positive Yamabe invariant $Y(M)$, the following functional is an upper bound of $R_{min}$ on $\mathcal{M}$:
@@
Moreover, if a smooth metric $g_{0}$ attains equality, then it is a Yamabe metric (i.e., $Y(M,[g_{0}])=Y(M)$) and an Einstein metric. However, when the Yamabe invariant is positive, this is not the case in general (see Remark 2.1 below or [21, Chapter 3]). A Yamabe metric is expected to be standard in some sense (cf. [29, Section 1]), but it remains a difficult problem to know how to actually obtain it as a limit of the sequence of solutions to the Yamabe problem, and in what sense it is standard (see also Subsection 6.3 below).
@@
    $\begin{split}&~~~~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,1}>0\\ &\left(\mathrm{resp.}~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n}>0\right)\end{split}$
@@
    $\begin{split}&~~~~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n}>0.\\ &\left(\mathrm{resp.}~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,1}>0\right)\end{split}$
@@
  • All the assumptions of the above Theorems 1.3, 1.4, 1.5 and 1.6 especially imply that $R_{g}$ is not a constant on $M$ or
@@
  • Interestingly, Dahl–Kröncke [8, 17] recently also discovered a relation between stability of Einstein metrics and certain type of scalar curvature rigidity. Since every Einstein metric does not satisfy any of the above assumptions (see Remark 1.2 above), our theorems above cannot be applied to Einstein metrics.
According to Listing’s work [20], we define four types of rigidities of metrics involving scalar curvature.
@@
Let $M$ be a smooth manifold. A metric $g_{0}$ on $M$ is type I scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
Let $M$ be a smooth manifold. A metric $g_{0}$ on $M$ is type II scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
Let $M$ be a smooth manifold. A metric $g_{0}$ on $M$ is type III scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
Let $M$ be a smooth manifold. A metric $g_{0}$ on $M$ is type IV scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
@@
    $\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}-R_{g}(x)\cdot\lambda_{Ric,1}(x)\leq 0$
@@ 2 Previous rigidity results for metrics with positive scalar curvature
Let $(M_{0}^{n},g_{0})~(n\geq 3)$ be an oriented closed Riemannian manifold with nonnegative curvature operator, positive Ricci curvature and non-vanishing Euler characteristic. Suppose that $(M^{n},g)$ is an oriented closed Riemannian manifold and $f:M\rightarrow M_{0}$ is a spin map of non-vanishing $\hat{A}-$degree $\mathrm{deg}_{\hat{A}}(f)\neq 0$ and $\mathrm{area}(f)\leq 1$. Then $R_{g}\geq R_{g_{0}}\circ f$ implies that $R_{g}=R_{g_{0}}\circ f$. If moreover, $\mathrm{Ric}_{g}>0$ and $2\,\mathrm{Ric}_{g}-R_{g}\cdot g<0$ on $M$, then $f:M\rightarrow M_{0}$ is a Riemannian submersion.
@@
Let $(M,g)$ be a compact connected oriented Riemanniam manifold with nonnegatuve curvature operator on $\Lambda^{2}TM$, such that the universal covering of $M$ is homeomorphic to a symmetric space $G/K$ of compact type with $\mathrm{rk}G\leq\mathrm{rk}K+1$. Then $g$ is (globally) area-extremal. If moreover, $\mathrm{Ric}_{g}>0$ and $2\,\mathrm{Ric}_{g}-R_{g}\cdot g<0$ on $M$, then $R_{h}\geq R_{g}$ and $h\geq g$ on $\Lambda^{2}TM$ implies $h=g$.
@@
For a fixed Riemannian metric $g_{0}\in\mathcal{M}(M)$, we define the following two functionals $F_{1,g_{0}}$ and $F_{2,g_{0}}$.
$\begin{split}F_{1,g_{0}}&:\mathcal{M}(M)\rightarrow\mathbb{R}~;~g\mapsto\max_{M}R_{g_{0}}\cdot\|g_{0}\|^{2}_{1,g},\\ F_{2,g_{0}}&:\mathcal{M}(M)\rightarrow\mathbb{R}~;~g\mapsto\max_{M}R_{g_{0}}\cdot\|g_{0}\|_{2,g}.\end{split}$
@@
Suppose there is a metric $g\in\mathcal{M}(M)$ such that $R_{min}(g)=\min_{M}R_{g}>F_{1,g_{0}}(g)$. Then, from Definition 1.1, there is a positive constant $c>0$ such that $g=c\cdot g_{0}$. Hence, $R_{g}=c^{-1}R_{g_{0}}\leq c^{-1}F_{1,g_{0}}(g_{0})=F_{1,g_{0}}(g)$ on $M$. This contradicts our supposition $R_{min}(g)=\min_{M}R_{g}>F_{1,g_{0}}(g)$. The proof for the corresponding statement to $F_{2,g_{0}}$ is similar. ∎
Moreover, each equality is attained by the scalar curvature rigid metric $g_{0}$ if $R_{g_{0}}$ is constant on $M$. On the other hand, Gromov [11] introduced the $K$-area of $M$ and gave an upper bound of $R_{min}$ on closed spin $n$-manifolds (“$K$-area inequality” in [11, $5\frac{1}{4}$]), which is expressed using the $K$-area and the dimension $n$. See also [21] for more detail.
As pointed out in [21, Section 3], when the Yamabe invariant, alias the sigma constant, $Y(M^{n})$ is non-positive, then
@@ 3 Preliminaries
where the term $g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h$ is a contraction of three copies of $g^{-1}$ (i.e., $g$ with raised indices) and two copies of $\bar{\nabla}h=\bar{\nabla}g$. And, the term $(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}*\mathrm{Ric}_{\bar{g}}$ is the trace of $\mathrm{Ric}_{\bar{g}}$ with respect to $\left((\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}\right)^{-1}$. Note that $\bar{g}+h$ is positive definite if $\|h\|_{\bar{g}}$ is small enough. All the proofs of these formulas are given in Section 7 below.
@@
instead of (15), a more subtle deformation is needed (see [5, Theorem 4]).
@@
$\Delta_{g}u=\lambda u$
for some positive constant $\lambda>0$. Then, for any $r>0$, there is a small $s>0$ (resp. $s<0$) depending only on $n,g,\lambda,K$ and $r$ such that $g_{s,u}:=(1+su)g$ is a smooth metric in $B_{r}(p)$ and that
@@
First, we prove the case of $s>0.$ From formula (14) and Li-Wang’s gradient estimate [19, Theorem 6.1], there is a constant $C>0$ depending on $n$ such that